# Balance Correlations, Agentic Zeros, and Networks: The Structure of 192 Years of War and Peace

### v1

@@ 2.4 Triad Signature
  • {P} :=assign:=:= the set of ($i,j$) unordered pairs designating all ties where $i$ has a positive tie with $j$;
  • {N} :=assign:=:= the set of ($i,j$) unordered pairs designating all ties where $i$ has a negative tie with $j$;
  • {Z} :=assign:=:= the set of ($i,j$) unordered pairs designating all ties where $i$ does not have a positive tie nor a negative tie with $j$, which we indicate as a neutral or “zero” tie.
We ignore self-ties ($i,i$) and hence these will not be in these three sets.
By convention, we represent these three sets in square matrix form as follows:
  • $\textbf{P}_{ij}=\begin{cases}1,&\text{if ($i,j$)}\ \in\{\textbf{P}\}\\ 0,&\text{otherwise}\end{cases}$
  • $\textbf{N}_{ij}=\begin{cases}1,&\text{if ($i,j$)}\ \in\{\textbf{N}\}\\ 0,&\text{otherwise}\end{cases}$
  • $\textbf{Z}_{ij}=\begin{cases}1,&\text{if ($i,j$)}\ \in\{\textbf{Z}\}\\ 0,&\text{otherwise.}\end{cases}$
These sets are, by definition, mutually exclusive and exhaustive in the network; thus, their sum as well as their union forms a complete and simple graph on the full set of nodes in $\mathcal{G}$. These three networks (P, N and Z) henceforth will be referred to as the set of primitive relations in $\mathcal{G}$.
@@ 3 $A\stackrel{{\scriptstyle s}}{{\text{---}}}X$ (Alter with X)
Table 4: Triad Signature Position Designation: The triad signature is a 3-tuple. Each edge, $q,r,$ and $s$, designates an undirected tie of type $p,n,$ or $z$ among the triple labeled Ego, Alter and X.
@@ 2.4.1 Identifying the 27 Triad Signatures
The 27 possible labeled configurations are provided in Figure 1. Positive ties (p) are represented in black; negative ties (n) are represented in dashed red lines; zero-ties (z) or neutral ties are blank, and are indicated with a letter “z” to emphasize they are treated as ties. Node labels refer to the Ego (E), Alter (A), and Context (X) roles in each triad. As we will discuss later on, at the network level some of these triads will be obfuscated. Nonetheless, in order to keep the theoretical stories clear, it is still important to retain all 27 triad signatures.
@@
Note: Each signature is given by the three letters, any combination of p, n and z, in angle brackets. A black solid line indicates a positive tie; a red dashed line indicates a negative tie; no line indicates a zero-tie (a neutral or non-tie). Node roles of Ego, Alter, and Contextual third party are represented by the letters E, A and X, respectively.
1
@@ 2.4.2 Triad Signature as a Predictive Statement
To illustrate this feature, we begin with a list of all eight classic Heiderian balance-theory statements, along with our triadic signature notation in Table (5). The core prediction in Heider’s model is that if a triple is unbalanced, then this creates psychological tension and discomfort. Thus, an unbalanced triple should not occur, or if it does occur, an attempt will be made by Ego to balance it by reconsidering her relations with Alter to make it balanced. On the other hand, if a triple is balanced, it will induce a sense of equanimity on Ego’s part and is more likely to be stable and remain in place.
@@
Table 5: Rapaport descriptions and their corresponding Triad Signature representations.
Viewed in this way, each balanced signature easily translates to a corresponding prediction from balance theory. Consider the following decomposition of the triadic elements.
@@
As a general principle, if the triple <q.rs> is balanced, then the prediction is:
$P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X)\quad>\quad P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X))\hss$ (7)
To illustrate, the Heiderian balance prediction that “the enemy of an enemy is a friend” can be restated more formally as “the probability that a specific Alter will be a friend is enhanced if I have an enemy who is also the enemy of this Alter.” This statement immediately corresponds to and is represented by the configuration <p.nn>, where p is the dependent variable and $nn$ is the 2-path conditional. The prediction is that $p$ is more likely to occur given the presence of the 2-path $nn$ than if $nn$ is not present.
$P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X)\quad<\quad P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X))\hss$ (8)
@@ 2.4.3 Triad Signature-induced Measure of Network Balance: The Balance Correlation
The Embeddedness Principle. Before we introduce this third, and most critical, function of the Triad Signature, we note one final difference between Heider’s original theoretical formulation and applications to social network analysis. Heider himself focused on isolated and ideational triples of entities and made predictions based solely on these three entities. In a network context, however, it is important to recognize that no labeled Ego  X  Alter combination occurs in isolation but rather is embedded in a complex structure of triples. In fact, changes in one tie can imply balance changes elsewhere in the network. Consider the case depicted in Figure 2. The three nodes in the middle of the graph (Ego, 1, and 2), constitute an unbalanced triple: Ego has a positive tie with Node 1, and Ego also has a positive tie with Node 2; but, Nodes 1 and 2 have a negative tie. Consider the specific ordered triple where node 1 is Alter and node 2 is the contextual X. Ego is faced with a dilemma. In considering whether to change her relationship with Alter (node 1), Ego finds herself embedded in several critical triples with node 1 as an Alter ((Ego,1,3), (Ego,1,4), (Ego,1,5)). Each of these triples has the signature of <p.pp>, and therefore these are balanced. If Ego were to balance the (Ego,1,2) triple by switching her relationship with node 1 from <p> to <n>, then that would distort the balance that currently exists in these three other <p.pp> triples. In Heiderian terms, the balance cost would far outweigh the balance benefits for switching.
@@
Note: The black edges represent positive ties; the red dashed edges represent negative ties. The focal labeled triple of interest here is {Ego,1,2}, which is unbalanced (<p.pn>). Since the Ego-2 tie is embedded in more unbalanced triples than the Ego-1 tie is, Ego is more likely to change her relationship with node 2 as an Alter than node 1 as an Alter in trying to resolve this unbalance.
@@
Taking this Embeddedness Principle into account, we now make Heider’s prediction more concrete. Since social networks provide many observations, even around a specific dyad, this affords us the opportunity to turn the prediction into a measurable probabilistic statement, rather than simply “the network is balanced” or “the network is unbalanced”, as Cartwright and Harary, (1956) originally suggested. We propose that balance statements are better framed with words like “tends to” or “is more likely to”. Consider the “friend of a friend is a friend” prediction. Rather than stating “In every case, a friend of a friend is a friend”, we argue that a better, empirically testable extension of this claim would be “A friend of a friend will tend to be a friend.” And the more friend-of-friend contexts one is embedded in, the more likely Ego will choose Alter as a friend, assuming balance theory is correct. This insight will help us explore Heider’s model more precisely.
@@
One of the basic predictions stemming from balance theory is that, if individuals are prone to attaining and maintaining balance, then at any given time we should find that balanced relations are more prevalent in a network. That is, if balance theory actually works to influence overall structures, then we should find that the 2-paths that exist in a network should be associated with a higher probability of the balance-confirming direct tie from Ego to Alter. As we observed in equation 1, the more a balance-based 2-path conditional exists around any particular Ego  Alter pair, the higher the probability we would observe the predicted direct tie from Ego to Alter.
@@
A key advantage of formalizing Heiderian predictions using triad signatures is that they make it easy to extend this correlational measure to include this embeddedness principle. Dekker, et al. (2019) created a correlation coefficient, the Transitivity Correlation, that assesses the extent to which the 2-path pre-transitive condition in a digraph (${i}\text{-}{k}\text{-}{j}$) is correlated with the presence of a direct tie ${i}\text{-}{j}$. Their solution to this problem was simply to correlate the off-diagonal elements in the matrix form of $\mathcal{G}$ with the corresponding elements in matrix $\mathcal{G}^{2}$. Since $\mathcal{G}^{2}$ contains the number of directed 2-path ties of the form $i\text{-}k\text{-}j$ for every $\langle i,j\rangle$ dyad, this point biserial correlation indicates the extent to which there is an overall tendency in the digraph $\mathcal{G}$ to close a directed 2-path with a directed tie. If the correlation is strong and positive, that would indicate that the tendency towards transitive closure is increased as a function of the number of directed 2-path paths surrounding any dyad. A negative correlation would indicate that the more directed 2-path paths that surround a dyad, the less likely it is that $i$ will be directly tied to $j$, indicating that $\mathcal{G}$ has an intransitivity propensity (such as one would find in a star-shaped graph). A zero correlation would indicate that, controlling for the density in $\mathcal{G}$ and the number of directed 2-paths in $\mathcal{G}$, there is no overall propensity towards transitivity in $\mathcal{G}$.
@@
${\textbf{M}_{1}}=\textbf{ Q},\ {\textbf{M}_{2}}=\textbf{RS}$
@@
A balance correlation, then, is simply the product-moment correlation between off-diagonal elements in $\textbf{M}_{1}$ with the corresponding elements in $\textbf{M}_{2}$. Moreover, as just demonstrated, these correlations are easily generated from the signatures themselves. For example, the <n.pn> signature generates a balance statement that negative n-ties are induced from an abundance of 2-path pn-ties; and the balance correlation is $\rho(\bf N,PN)$, where N is the matrix of n-ties, and PN is the compound word (White, Boorman & Breiger, 1976) created by the matrix product of the P matrix and the N matrix of ties .
@@
Phenotypical Twins. As we have already noted above, there are $273^{3}=$ 27 triad signatures. And, each of these signatures doubles as a possible prediction and generates a correlation that permits us to evaluate this prediction. It should be recognized, while these predictive statements and measured correlations can be generated for any of the 27 signatures, many of these signatures do not correspond to anything that Heider would claim actually stems from balance theory. Yet, it is still possible that these empirical correlations might yield insights into structural patterns that could lead to new theoretical extensions, akin to what balance theory has produced over the past half century (e.g., Holland and Leinhardt,, 1979; Davis,, 1979). A thorough exploration of such patterns in any dataset would result in a 9 $\times$ 3 correlation matrix, where each of the three dependent relations (Q, R, and S) is potentially predicted by each of the 9 possible 2-path relations.
@@
We provide the following theorem for uncovering twins in triad signatures (a general proof of this theorem on multi-digraphs is provided in the Appendix):
@@
Thus, <n.pn> and <n.np> are twins. Note, however, that <n.pn> and <p.nn> are not twins, even though they are isomorphic in their graph structures  [^1] It is important to note that non-twin signatures can, in any given graph, have identical balance correlations. There may be several balance correlations equal to 0, for example, indicating that there is no tendency toward balance or unbalance as predicted in those signatures. The critical issue with twin signatures is that their Balance correlations will be equal by construction in any and all possible graphs of any size composed of any arrangement of ¡p¿, ¡n¿ and ¡z¿ ties.. As we will see below in our analysis, isomorphic triads but which are not twin signatures, can have dramatically different balance correlations. Indeed, in some cases, we will observe isomorphic triadic signatures that will reveal opposite signs; that is, one signature will exhibit balance while its isomorphic counterpart will be unbalanced.

### v2

@@ 2.4 Triad signature
$\{\textbf{P}\}:=$
the set of ($i,j$) unordered pairs designating all ties where $i$ has a positive tie with $j$;
$\{\textbf{N}\}:=$
the set of ($i,j$) unordered pairs designating all ties where $i$ has a negative tie with $j$; and
$\{\textbf{Z}\}:=$
the set of ($i,j$) unordered pairs designating all ties where $i$ does not have a positive tie nor a negative tie with $j$, which we indicate as a neutral or “zero” tie.
We ignore self-ties $(i,i)$ and hence these will not be in these three sets. By convention, we represent these three sets in square matrix form as follows:
$\textbf{P}_{ij}=\begin{cases}1&\text{if ($i,j$)}\ \in\{\textbf{P}\}\\ 0&\text{otherwise},\end{cases}\quad\textbf{N}_{ij}=\begin{cases}1&\text{if ($i,j$)}\ \in\{\textbf{N}\}\\ 0&\text{otherwise},\end{cases}\quad\textbf{Z}_{ij}=\begin{cases}1&\text{if ($i,j$)}\ \in\{\textbf{Z}\}\\ 0&\text{otherwise.}\end{cases}$
These sets are, by definition, mutually exclusive and exhaustive in the network; thus, their sum, as well as their union, forms a complete and simple graph on the full set of nodes in $\mathcal{G}$. These three networks (P, N and Z) henceforth will be referred to as the set of primitive relations in $\mathcal{G}$.
@@ 3 $A\stackrel{{\scriptstyle s}}{{\text{---}}}X$ (Alter with X)
Note: The triad signature is a 3-tuple. Each edge, $q,r,$ and $s$, designates an undirected tie of type $p,n,$ or $z$ among the triple labeled Ego, Alter and X.
@@ 2.4.1 Identifying the 27 triad signatures
The 27 possible labeled configurations are provided in Fig 1. Positive ties (p) are represented in black; negative ties (n) are represented in dashed red lines; zero-ties (z) or neutral ties are blank, and are indicated with a letter “z” to emphasize they are treated as ties. Node labels refer to the Ego (E), Alter (A), and Context (X) roles in each triad. As we will discuss later, at the network level some of these triads will be mathematically redundant. Nonetheless, in order to keep the theoretical stories clear, it is still important to retain all 27 triad signatures.
@@
EAXppp<$p.pp$>
EAXnpp<$n.pp$>
EAXzpp<$z.pp$>
EAXppn<$p.pn$>
EAXnpn<$n.pn$>
EAXzpn<$z.pn$>
EAXppz<$p.pz$>
EAXnpz<$n.pz$>
EAXzpz<$z.pz$>
EAXpnp<$p.np$>
EAXnnp<$n.np$>
EAXznp<$z.np$>
EAXpnn<$p.nn$>
EAXnnn<$n.nn$>
EAXznn<$z.nn$>
EAXpnz<$p.nz$>
EAXnnz<$n.nz$>
EAXznz<$z.nz$>
EAXpzp<$p.zp$>
EAXnzp<$n.zp$>
EAXzzp<$z.zp$>
EAXpzn<$p.zn$>
EAXnzn<$n.zn$>
EAXzzn<$z.zn$>
EAXpzz<$p.zz$>
EAXnzz<$n.zz$>
EAXzzz<$z.zz$>
Note: Each signature is given by any combination of p, n and z, in angle brackets. A solid line indicates a positive tie; a dashed line indicates a negative tie; no line indicates a zero-tie. Node labels indicate roles of Ego (E), Alter (A), and Contextual (X).
@@ 2.4.2 Triad signature as a predictive statement
To illustrate this feature, we begin with a list of all eight classic Heiderian balance-theory statements, along with our triadic signature notation in Table 5. The core prediction in Heider’s model is that if a triple is unbalanced, then this creates psychological tension and discomfort. Thus, an unbalanced triple should not occur, or if it does occur, an attempt will be made by Ego to balance it by reconsidering her relations with Alter to make it balanced. On the other hand, if a triple is balanced, it will induce a sense of equanimity on Ego’s part and is more likely to be stable and remain in place. Viewed in this way, each balanced signature easily translates to a corresponding prediction from balance theory. Consider the following decomposition of the triadic elements.
Table 5: Rapaport descriptions and the corresponding triad signature representations.
@@
As a general principle, if the triple <q.rs> is balanced, then the prediction is
$P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X)\quad>\quad P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X))\hss.$ (7)
To illustrate, the Heiderian balance prediction that “the enemy of an enemy is a friend” can be restated more formally as “the probability that a specific Alter will be a friend is enhanced if I have an enemy who is also the enemy of this Alter.” This statement immediately corresponds to and is represented by the configuration <p.nn>, where p is the dependent variable and $nn$ is the 2-path conditional. The prediction is that $p$ is more likely to occur given the presence of the 2-path $nn$ than if $nn$ is not present:
$P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X)\quad<\quad P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X))\hss.$ (8)
@@ 2.4.3 Triad signature-induced measure of network balance: The balance correlation
The Embeddedness Principle. Before we introduce this third, and most critical, function of the Triad Signature, we note one final difference between Heider’s original theoretical formulation and applications to social network analysis. Heider himself focused on isolated and ideational triples of entities and made predictions based solely on these three entities. In a network context, however, it is important to recognize that no labeled Ego  X  Alter combination occurs in isolation but rather is embedded in a complex structure of triples. In fact, changes in one tie can imply balance changes elsewhere in the network. Consider the case depicted in Fig  2. The three nodes in the middle of the graph (Ego, 1, and 2), constitute an unbalanced triple: Ego has a positive tie with Node 1, and Ego also has a positive tie with Node 2; but, Nodes 1 and 2 have a negative tie. Consider the specific ordered triple where node 1 is Alter and node 2 is the contextual X. Ego is faced with a dilemma. In considering whether to change her relationship with Alter (node 1), Ego finds herself embedded in several critical triples with node 1 as an Alter ((Ego,1,3), (Ego,1,4), (Ego,1,5)). Each of these triples has the signature of <p.pp>, and therefore these are balanced. If Ego were to balance the (Ego,1,2) triple by switching her relationship with node 1 from <p> to <n>, then that would distort the balance that currently exists in these three other <p.pp> triples. In Heiderian terms, the balance cost would far outweigh the balance benefits for switching.
@@
Ego12345678
Note: The black edges represent positive ties; the dashed edges represent negative ties. The focal labeled triple of interest here is {Ego,1,2}, which is unbalanced (<p.pn>). Since the Ego-2 tie is embedded in more unbalanced triples than the Ego-1 tie is, Ego is more likely to change her relationship with node 2 as an Alter than node 1 as an Alter in trying to resolve this unbalance.
@@
Taking this Embeddedness Principle into account, we now make Heider’s prediction more concrete. Since social networks provide many observations, even around a specific dyad, this affords us the opportunity to turn the prediction into a measurable probabilistic statement, rather than simply “the network is balanced” or “the network is unbalanced”, as [6] originally suggested. We propose that balance statements are better framed with words like “tends to” or “is more likely to”. Consider the “friend of a friend is a friend” prediction. Rather than stating “In every case, a friend of a friend is a friend”, we argue that a better, empirically testable extension of this claim would be “A friend of a friend will tend to be a friend.” And the more friend-of-friend contexts one is embedded in, the more likely Ego will choose Alter as a friend, assuming balance theory is correct. This insight will help us explore Heider’s model more precisely.
@@
One of the basic predictions stemming from balance theory is that, if individuals are prone to attaining and maintaining balance, then at any given time we should find that balanced relations are more prevalent in a network. That is, if balance theory actually works to influence overall structures, then we should find that the 2-paths that exist in a network should be associated with a higher probability of the balance-confirming direct tie from Ego to Alter. As we observed in Eq 1, the more a balance-based 2-path conditional exists around any particular Ego  Alter pair, the higher the probability we would observe the predicted direct tie from Ego to Alter.
@@
A key advantage of formalizing Heiderian predictions using triad signatures is that they make it easy to extend this correlational measure to include this embeddedness principle. [8] created a correlation coefficient, the Transitivity Correlation, that assesses the extent to which the 2-path pre-transitive condition in a digraph (${i}\text{-}{k}\text{-}{j}$) is correlated with the presence of a direct tie ${i}\text{-}{j}$. Their solution to this problem was simply to correlate the off-diagonal elements in the matrix form of $\mathcal{G}$ with the corresponding elements in matrix $\mathcal{G}^{2}$. Since $\mathcal{G}^{2}$ contains the number of directed 2-path ties of the form $i\text{-}k\text{-}j$ for every $\langle i,j\rangle$ dyad, this point biserial correlation indicates the extent to which there is an overall tendency in the digraph $\mathcal{G}$ to close a directed 2-path with a directed tie. If the correlation is strong and positive, that would indicate that the tendency towards transitive closure is increased as a function of the number of directed 2-path paths surrounding any dyad. A negative correlation would indicate that the more directed 2-path paths that surround a dyad, the less likely it is that $i$ will be directly tied to $j$, indicating that $\mathcal{G}$ has an intransitivity propensity (such as one would find in a star-shaped graph). A zero correlation would indicate that, controlling for the density in $\mathcal{G}$ and the number of directed 2-paths in $\mathcal{G}$, there is no overall propensity towards transitivity in $\mathcal{G}$.
@@
${\textbf{M}_{1}}=\textbf{ Q},\ {\textbf{M}_{2}}=\textbf{RS},$
@@
A balance correlation, then, is simply the product-moment correlation between off-diagonal elements in $\textbf{M}_{1}$ with the corresponding elements in $\textbf{M}_{2}$. Moreover, as just demonstrated, these correlations are easily generated from the signatures themselves. For example, the <n.pn> signature generates a balance statement that negative n-ties are induced from an abundance of 2-path pn-ties; and the balance correlation is $\rho(\bf N,PN)$, where N is the matrix of n-ties, and PN is the compound word [46] created by the matrix product of the P matrix and the N matrix of ties .
@@
Phenotypical Twins. As we have already noted above, there are $273^{3}=$ 27 triad signatures. And, each of these signatures doubles as a possible prediction and generates a correlation that permits us to evaluate this prediction. It should be recognized, while these predictive statements and measured correlations can be generated for any of the 27 signatures, many of these signatures do not correspond to anything that Heider would claim actually stems from balance theory. Yet, it is still possible that these empirical correlations might yield insights into structural patterns that could lead to new theoretical extensions, akin to what balance theory has produced over the past half century [27, 7]. A thorough exploration of such patterns in any dataset would result in a 9 $\times$ 3 correlation matrix, where each of the three dependent relations (Q, R, and S) is potentially predicted by each of the 9 possible 2-path relations.
@@
We provide the following theorem for uncovering twins in triad signatures (a general proof of this theorem on multi-digraphs is provided in S1 Appendix.):
@@
Thus, <n.pn> and <n.np> are twins. Note, however, that <n.pn> and <p.nn> are not twins, even though they are isomorphic in their graph structures. It is important to note that non-twin signatures can, in any given graph, have identical balance correlations. There may be several balance correlations equal to 00, for example, indicating that there is no tendency toward balance or unbalance as predicted in those signatures. The critical issue with twin signatures is that their Balance correlations will be equal by construction in any and all possible graphs of any size composed of any arrangement of <p>, <n> and <z> ties. As we will see below in our analysis, isomorphic triads but which are not twin signatures, can have dramatically different balance correlations. Indeed, in some cases, we will observe isomorphic triadic signatures that will reveal opposite signs; that is, one signature will exhibit balance while its isomorphic counterpart will be unbalanced.